Calculate the specific heat capacity of titanium if a 45 g sample absorbs 476 J as its temperature changes from 20⁰C to 40.25⁰C. _ J/g ⁰C (2 decimals)
1
Expert's answer
2019-05-16T06:25:08-0400

Solution.

"Q = c \times m \times \Delta t"

"Q = c \times m \times (t2-t1)"

"c = \frac{Q}{m \times \Delta t}"

"c = \frac{476}{45 \times (40.25-20)}"

c = 0.52 J/g ⁰C

Answer:

c = 0.52 J/g ⁰C